Two charges of +8.0mC and -6.0 mC attract each other with a force of 3.0x103 N in a vacuum. What is the distance between the charges.

1
Expert's answer
2022-02-13T13:01:45-0500

Given:

"q_1=\rm 8.0\:mC=8.0*10^{-3}\:\rm C"

"q_2=\rm 6.0\:mC=6.0*10^{-3}\:\rm C"

"F=\rm 3.0*10^{3}\:\rm N"


The Coulomd law says

"F=k\frac{q_1q_2}{r^2}"

Hence, the distance between the charges

"r=\sqrt{\frac{kq_1q_2}{F}}\\
=\sqrt{\frac{9*10^9*8.0*10^{-3}*6.0*10^{-3}}{3.0*10^{3}}}=12\;\rm m"